# Learning to Evolve with Convergence Guarantee via Neural Unrolling

### v1

@@ V-A Experimental Setting
Benchmark Setup. We evaluate the optimization efficiency and zero-shot generalization capability of L2E using a hierarchical benchmark suite. This suite ranges from synthetic functions to real-world robot control tasks. We structure the evaluation protocol into three categories based on the relationship between the training and testing distributions:
  • Synthetic Benchmarks (In-Distribution): We utilize BBOB-10D, BBOB-30D [54], and LSGO-1000D [55] as the primary testbeds. These datasets constitute both the training sets and the in-distribution test sets. Specifically, the BBOB tasks evaluate basic optimization capabilities with budgets of $2\times 10^{4}$ (10D) and $5\times 10^{4}$ (30D). In contrast, LSGO assesses scalability in high-dimensional regimes ($D\approx 1000$) with a budget of $3\times 10^{6}$.
  • Out-of-Distribution (OOD) Generalization: We select two benchmarks to assess the transferability of L2E across disparate domains. First, we employ the UAV benchmark [56], which comprises 56 real-world path planning instances. Here, we test a model trained exclusively on BBOB-30D under a constrained budget of $2.5\times 10^{2}$. Second, we utilize BBOB-surrogate-10D, i.e., a synthetic variant derived from surrogate model landscapes. This benchmark tests robustness against landscape shifts and approximation errors. In this case, we apply the model trained on standard BBOB-10D directly without fine-tuning.
  • Neuroevolution for Robot Control (Flexibility Verification): We extend the evaluation to the NE benchmark [57], which consists of high-dimensional robot control policy search tasks ($D>1000$). We use this suite specifically to benchmark L2E against reinforcement learning methods. This comparison assesses the flexibility of the framework in control scenarios, given a maximum evaluation budget of $2,500$.
@@
Method Sharp_Rid Different_Pow Schaffers_hig Composite_Gri Schwefel Gallagher_21P Katsuura Lunacek_bi_Ras
PSO 9.431E+02 7.173E+00 2.411E+01 4.266E+00 3.235E+02 6.513E+01 1.209E+00 1.166E+02
($\pm$1.627E+02) ($\pm$3.197E+00) ($\pm$9.562E-01) ($\pm$6.784E-01) ($\pm$3.334E+02) ($\pm$5.339E+00) ($\pm$2.412E-01) ($\pm$7.371E+00)
DE 2.583E+02 2.088E+00 1.081E+01 4.509E+00 2.413E+00 1.420E+01 2.733E+00 8.125E+01
($\pm$2.861E+01) ($\pm$1.005E+00) ($\pm$8.182E-01) ($\pm$2.996E-01) ($\pm$9.219E-02) ($\pm$1.058E+01) ($\pm$7.799E-01) ($\pm$3.929E+00)
CMAES 1.309E+03 1.309E+01 4.228E+01 3.101E+00 1.671E+03 7.707E+01 5.306E-01 6.387E+01
($\pm$1.568E+00) ($\pm$4.955E-02) ($\pm$5.127E-03) ($\pm$5.885E-01) ($\pm$4.225E+01) ($\pm$1.659E-02) ($\pm$6.150E-02) ($\pm$4.391E-02)
SAHLPSO 3.679E+02 1.316E+00 1.229E+01 5.759E+00 3.942E+01 1.675E+01 1.577E+00 9.341E+01
($\pm$2.165E+01) ($\pm$1.267E+00) ($\pm$5.473E+00) ($\pm$5.528E-01) ($\pm$3.132E+01) ($\pm$5.020E+00) ($\pm$5.575E-01) ($\pm$2.748E+00)
RNNOPT 1.798E+03 1.069E+01 5.223E+01 2.056E+00 8.060E+03 8.486E+01 3.440E+00 1.221E+02
DEDQN 1.057E+03 1.047E+01 3.073E+01 1.127E+01 4.618E+03 5.834E+01 3.613E+00 1.480E+02
($\pm$1.827E+02) ($\pm$3.672E+00) ($\pm$4.258E+00) ($\pm$1.689E+00) ($\pm$8.773E+02) ($\pm$9.502E+00) ($\pm$5.164E-01) ($\pm$2.998E+01)
LES 3.274E+03 1.505E+03 7.211E+02 1.700E+03 2.104E+03 2.454E+03 6.017E+02 1.270E+03
($\pm$6.658E+01) ($\pm$3.239E-01) ($\pm$6.131E-01) ($\pm$3.196E-02) ($\pm$2.379E-01) ($\pm$1.005E+01) ($\pm$1.436E-01) ($\pm$2.917E+00)
LGA 9.683E+02 1.663E+01 2.613E+01 1.322E+01 1.234E+04 4.875E+01 1.618E+00 1.778E+02
($\pm$2.034E+02) ($\pm$5.850E+00) ($\pm$4.232E+00) ($\pm$5.755E+00) ($\pm$3.810E+03) ($\pm$1.666E+01) ($\pm$0.000E+00) ($\pm$1.977E+01)
GLHF 1.111E+03 1.530E+01 2.792E+01 9.223E+00 4.686E+03 4.876E+01 3.299E+00 1.499E+02
($\pm$2.052E+02) ($\pm$8.705E-01) ($\pm$3.296E+00) ($\pm$1.221E+00) ($\pm$1.331E+03) ($\pm$1.684E+01) ($\pm$4.095E-01) ($\pm$2.572E+01)
B2OPT 4.721E+02 4.665E+00 1.176E+01 1.633E+00 2.877E+00 2.987E+01 2.293E+00 7.342E+01
($\pm$6.671E+01) ($\pm$1.035E+00) ($\pm$2.675E+00) ($\pm$1.135E-01) ($\pm$3.331E-01) ($\pm$1.920E+01) ($\pm$4.472E-02) ($\pm$5.952E+00)
[…]
indicates ours variant enhanced with the proposed Mamba-based neural operator.
@@
Method Sharp_Rid Different_Pow Schaffers_hig Composite_Gri Schwefel Gallagher_21P Katsuura Lunacek_bi_Ras
PSO 1.360E+03 1.928E+01 2.751E+01 5.665E+00 2.016E+01 6.659E+01 3.208E+00 3.461E+02
($\pm$1.372E+02) ($\pm$3.804E+00) ($\pm$4.309E+00) ($\pm$2.647E-01) ($\pm$1.457E+01) ($\pm$2.572E+00) ($\pm$2.201E-01) ($\pm$2.614E+01)
DE 7.012E+02 8.848E+00 1.536E+01 6.595E+00 2.439E+02 1.083E+01 3.000E+00 3.043E+02
($\pm$3.727E+01) ($\pm$7.387E-01) ($\pm$5.156E+00) ($\pm$1.962E-01) ($\pm$1.181E+02) ($\pm$7.941E-01) ($\pm$1.823E-01) ($\pm$3.679E+01)
CMAES 1.485E+03 2.591E+01 4.616E+01 4.878E+00 3.021E+00 7.824E+01 4.913E-01 2.853E+02
($\pm$1.474E+00) ($\pm$9.453E-02) ($\pm$4.374E-02) ($\pm$1.767E-01) ($\pm$6.518E-02) ($\pm$7.518E-02) ($\pm$1.826E-01) ($\pm$1.699E+00)
SAHLPSO 1.500E+03 1.707E+01 3.268E+01 8.649E+00 4.673E+03 4.442E+01 3.192E+00 4.744E+02
($\pm$3.082E+01) ($\pm$8.338E+00) ($\pm$4.166E+00) ($\pm$1.578E+00) ($\pm$1.368E+03) ($\pm$8.733E+00) ($\pm$7.748E-01) ($\pm$2.208E+01)
RNNOPT 2.314E+03 7.363E+01 7.426E+01 6.430E+00 2.606E+04 8.447E+01 7.372E+00 4.541E+02
DEDQN 2.360E+03 4.823E+01 4.896E+01 2.184E+01 6.849E+04 7.642E+01 4.808E+00 8.033E+02
($\pm$1.659E+02) ($\pm$2.662E+00) ($\pm$5.639E+00) ($\pm$1.482E+00) ($\pm$1.161E+04) ($\pm$3.011E+00) ($\pm$3.796E-01) ($\pm$9.052E+01)
LES 2.895E+03 1.310E+02 1.341E+03 1.300E+03 7.007E+03 1.180E+03 1.303E+03 1.749E+03
($\pm$3.334E+01) ($\pm$1.310E+00) ($\pm$2.615E+00) ($\pm$4.597E-04) ($\pm$1.338E+03) ($\pm$1.277E+00) ($\pm$2.324E-01) ($\pm$1.423E+01)
LGA 2.821E+03 9.115E+01 6.749E+01 2.159E+01 9.618E+04 8.395E+01 3.231E+00 9.121E+02
($\pm$1.272E+02) ($\pm$2.051E+01) ($\pm$5.473E+00) ($\pm$3.009E+00) ($\pm$1.232E+04) ($\pm$1.719E+00) ($\pm$2.440E-01) ($\pm$1.066E+02)
GLHF 2.611E+03 6.510E+01 5.337E+01 2.541E+01 7.450E+04 8.305E+01 6.515E+00 8.946E+02
($\pm$1.152E+02) ($\pm$5.242E+00) ($\pm$5.011E+00) ($\pm$1.609E+00) ($\pm$4.114E+03) ($\pm$1.280E+00) ($\pm$1.767E-01) ($\pm$3.887E+01)
B2OPT 1.393E+03 1.921E+01 3.485E+01 2.156E+00 1.671E+03 7.715E+01 4.219E+00 3.267E+02
($\pm$1.978E+02) ($\pm$3.729E-01) ($\pm$1.291E+00) ($\pm$2.386E-01) ($\pm$8.537E+02) ($\pm$3.433E+00) ($\pm$2.024E-01) ($\pm$5.097E+00)
[…]
@@
indicates ours variant enhanced with the proposed Mamba-based neural operator.
@@
Compared Methods. We conduct a comparative analysis against representative algorithms to evaluate the effectiveness of the proposed L2E framework. This set encompasses both classical heuristics and learning-based optimizers. To situate our non-expansive unrolling paradigm within the broader L2O landscape, we incorporate two categories of learned algorithms. The first category consists of meta-learning-based neuro-evolution strategies, i.e., DEDQN [58], LES [28], and LGA [59]. The second category includes control-based solution generation models, i.e., RNNOPT [60], GLHF [61], and B2OPT [23]. Additionally, we employ established traditional algorithms, i.e., CMA-ES [11], DE [62], PSO [31], and SAHLPSO [63], which serve as performance anchors. We re-evaluate all competing methods within a unified experimental protocol and identical computational budgets, thereby ensuring a consistent assessment. Finally, we conduct all experiments on a single NVIDIA RTX A6000 GPU (PyTorch 2.0.1) and evaluate each task over 10 independent random seeds to report the mean and standard deviation.
V-B Performance on Numerical Benchmarks
We first examine the optimization capability of L2E on numerical landscapes where the meta-training and testing distributions align. This assessment covers two critical dimensions: precision on standard numerical problems (BBOB) and scalability to high dimensions (LSGO).
V-B1 Precision and Stability on Standard Benchmarks
Table I and Table II report the comparative results on BBOB-10D and BBOB-30D, respectively. L2E demonstrates a substantial advantage over state-of-the-art methods:
  • Performance Superiority and Architectural Efficacy: L2E establishes a consistent performance advantage. Specifically, it achieves the lowest mean error on 31 out of 32 test instances. On ill-conditioned and multimodal functions (e.g., Rosenbrock_rot, Lunacek_bi_Ras), L2E substantially outperforms strong baselines, such as GLHF and B2OPT. We attribute this escape capability to the non-expansive unrolling mechanism, which enforces a strict contractive constraint to prevent divergence in high-curvature valleys. Furthermore, the comparison between variants reveals that the Mamba-enhanced model (Ours) consistently improves upon the base model (Ours), particularly on non-separable landscapes (e.g., Ellipsoidal_hc). This result confirms that the structured state-space model offers a superior inductive bias for capturing complex variable couplings, in contrast to the generic attention or coordinate-wise mechanisms used in the standard version.
  • Convergence Trajectory Analysis: Figure 4 and Figure 5 illustrate the convergence behavior. L2E (red curves) exhibits a steeper initial descent and tighter confidence intervals. On unimodal functions (Bent Cigar), the model rapidly saturates to the optimum. This behavior demonstrates the effective exploitation of gradient-like information. Conversely, on multimodal tasks (Gallagher 21Peaks), L2E maintains steady progress without early stagnation. This balance confirms the efficacy of the Elite-Mask Fusion, which adaptively arbitrates between global exploration and local gradient refinement.
V-B2 Scalability to High-Dimensional Regimes
The results on the 1000D LSGO benchmark (Table III) further highlight the structural advantages of L2E. Specifically, L2E ranks first across all six representative high-dimensional functions. On the complex Shifted_Rastrigin, the model achieves a lower objective value, a result that surpasses LGA and GLHF. Existing Transformer-based optimizers (e.g., GLHF) encounter substantial bottlenecks in 1000D regimes due to the quadratic complexity of the self-attention mechanism. In contrast, our Mamba-based operator maintains linear complexity, which facilitates stable gradient propagation during meta-training. The ECDF curves in Fig. 6 and boxplots in Fig. 7 illustrate this distinction. While RNNOPT and DEDQN exhibit flatlining trajectories, which indicate gradient vanishing or saturation, L2E maintains a continuous descent. This behavior demonstrates the capability of L2E to navigate high-dimensional manifolds efficiently.
@@ V-B2 Scalability to High-Dimensional Regimes
Figure 4: Log-scaled convergence trajectories of our method and representative methods on the first 8 benchmark tasks from the BBOB-10D suite [54]. The x-axis denotes the number of function evaluations (FEs), while the y-axis reports the logarithm of the objective value (log cost), where lower values indicate superior performance. For clarity, only 8 representative tasks are visualized; Complete results on all 16 functions are provided in the supplementary material (i.e., Fig S1).
V-C Generalization to Real-World UAV Planning
A core contribution of L2E is its capability to generalize to unseen, real-world tasks without fine-tuning. We validate this Sim-to-Real transfer capability using the UAV Path Planning benchmark. Although the model is trained exclusively on synthetic BBOB-30D functions, L2E generalizes robustly to the 56 distinct UAV planning tasks (30D). As depicted in the radar plots of Fig. 8, L2E achieves the highest normalized performance across all seven terrain groups. The improvement margins over the second-best baseline range from 3.4% to 14.2%. This performance advantage stems from the proposed transfer mechanism. In contrast to reinforcement learning-based methods (e.g., DEDQN), which risk overfitting to specific landscape features encountered during training, L2E learns a generic descent operator that is governed by fixed-point properties. Consequently, the optimizer perceives the irregular constraints and noise profiles of the UAV simulator as intrinsic properties of the black-box function. This perception allows it to apply the learned contractive rules to navigate feasible regions.
V-D Neuroevolution for Robot Control
Finally, we assess the applicability of the framework to high-dimensional policy search tasks in the Neuroevolution (NE) benchmark [57]. This assessment represents a substantial domain shift from function optimization to control policy learning. The benchmark consists of six realistic robotic tasks across two domains, i.e., ant and humanoid. Specifically, each domain contains three complexity variants (ant-3, ant-4, ant-6; and humanoid-3, humanoid-4, humanoid-5). We benchmark the proposed method against GLEET [64], i.e., a specialized baseline that formulates neuroevolution within a reinforcement learning framework. GLEET trains an actor-critic agent to evolve population-level solutions via environmental interactions. We conduct this comparison to demonstrate that the learned optimizer can replace heuristic operators to optimize actor-critic evolutionary control policies.
Fig. 9 presents the comparative results, utilizing a normalized metric to enable intuitive visual comparison across tasks with distinct reward scales. The vertical axis represents deviations from a task-specific mean, rescaled by a common exponential factor: $y_{\text{plot}}=\frac{y-\text{mean}}{10^{\text{exp}}}.$ This transformation highlights the convergence dynamics around a central baseline, and it preserves cross-task comparability. On representative mid-difficulty tasks, such as ant-4 and humanoid-4, L2E exhibits steeper convergence trajectories and lower terminal costs than the baseline. Furthermore, the aggregated boxplots (Right) reveal that L2E achieves lower final objective values with consistently reduced variance across all six tasks. This result suggests that replacing heuristic population updates with the learned non-expansive operator stabilizes the policy learning process, thereby mitigating the instability typically associated with high-variance reward signals in RL environments.
@@ V-D Neuroevolution for Robot Control
Figure 5: Log-scaled convergence trajectories of our method and representative methods on the first 8 benchmark tasks from the BBOB-30D suite [54]. The x-axis denotes the number of function evaluations (FEs), while the y-axis reports the logarithm of the objective value (log cost), where lower values indicate superior performance. For clarity, only 8 representative tasks are visualized; Complete results on all 16 functions are provided in the supplementary material (i.e., Fig S2).
@@
VI Ablation Analysis and Mechanistic Insights
To disentangle the contribution of each algorithmic component and justify our design choices, we conduct a systematic ablation study on the BBOB-10D suite. This analysis is structured into two parts: (1) an investigation of the neural parameterization strategy for the evolution operator, specifically Shared versus Unshared Weights; and (2) a dissection of the combined efficacy of the core modules, i.e., ProxyGrad, SoftGate, and Mamba.
VI-A Evolutionary Curriculum via Unshared Parameterization
A pivotal architectural decision in deep unrolling is the strategy for sharing the meta-parameters $\mathcal{O}_{\mathrm{evo}}(\cdot;\bm{\omega})$ of the evolution operator $\mathcal{O}_{\mathrm{evo}}$ across time steps $k=1\dots K$.
  • Shared ($\bm{\omega}_{\text{shared}}$): This strategy ties a single set of weights across all unrolled blocks, which forces the operator to learn a time-invariant update rule.
  • Unshared ($\bm{\omega}^{k}$): This strategy allocates independent weights for each block, which allows the operator to evolve distinct behaviors at different stages of the optimization process.

### v2

@@ V-A Experimental Setting
Benchmark Setup. We evaluate the optimization efficiency and zero-shot generalization capability of L2E using a hierarchical benchmark suite. This suite encompasses tasks ranging from synthetic functions to real-world robot control. Regarding hyperparameters, we set the relaxation factor $\alpha=0.9$, the gating temperature $\tau=1$, the unrolling horizon $K=200$, and the population size $B=100$, unless otherwise specified. For spectral normalization, we use the default coefficient $1.0$ for all linear layers. We conduct all experiments on a single NVIDIA RTX A6000 GPU (PyTorch 2.0.1). Each task is evaluated over 10 independent random seeds, and we report both the mean and standard deviation.
@@
indicates ours variant enhanced with the proposed Mamba-based neural operator. Due to space limitations, only a subset is shown here; the full table is provided in the Supplementary Material (Table S.I).
  • Synthetic Benchmarks (In-Distribution): We utilize BBOB-10D, BBOB-30D [20], and LSGO-1000D [33] as the primary testbeds. These benchmarks represent the same function families utilized during meta-training, although we conduct the evaluation on held-out runs or instances under identical protocols. Specifically, BBOB assesses general optimization capability with budgets of $2\times 10^{4}$ (10D) and $5\times 10^{4}$ (30D), whereas LSGO evaluates scalability in high-dimensional regimes ($D\approx 1000$) with a budget of $3\times 10^{6}$.
  • Out-of-Distribution (OOD) Evaluation: We incorporate two benchmarks to assess robustness beyond the training distribution. First, we employ the UAV benchmark [53], which consists of 56 real-world path planning instances. Specifically, we directly evaluate a model trained on BBOB-30D under a constrained budget of $2.5\times 10^{2}$ without fine-tuning. Second, we use BBOB-surrogate-10D, a synthetic variant derived from surrogate-induced landscapes, to test robustness to landscape shifts and approximation mismatch; we apply the model trained on BBOB-10D directly.
  • Neuroevolution for Robot Control: We additionally evaluate performance on the NE benchmark [22], which comprises high-dimensional policy search tasks ($D>1000$) for robot control. This suite assesses the applicability of L2E in control scenarios and benchmarks it against representative reinforcement learning methods under a maximum evaluation budget of 2,5002{,}500.
@@
indicates ours variant enhanced with the proposed Mamba-based neural operator. Due to space limitations, only a subset is shown here; the full table is provided in the Supplementary Material (Table S.II).
@@
Compared Methods. We conduct a comparative analysis against representative algorithms to evaluate the effectiveness of the proposed L2E framework, covering both classical heuristics and learning-based optimizers. To situate our stability-inducing unrolling framework within the broader L2O landscape, we include two categories of learned optimizers. The first category consists of meta-learning-based neuro-evolution strategies, including DEDQN [56], LES [27], and LGA [26]. The second category includes control-based solution generation models, including RNNOPT [59], GLHF [31], and B2OPT [32]. In addition, we include representative classical population-based baselines, including DE [54], PSO [25], and SAHLPSO [58], to provide reference performance under the same evaluation budgets. We re-evaluate all competing methods under a unified experimental protocol and identical computational budgets.
@@
Figure 4: Log-scaled convergence trajectories of our method and representative methods on the first 8 benchmark tasks from the BBOB-10D suite [20]. Complete results on all 16 functions are provided in the supplementary material (i.e., Fig S1).
V-B Performance on Numerical Benchmarks
We first examine the optimization capability of L2E on numerical landscapes where the meta-training and testing distributions align. This assessment focuses on two aspects: performance on standard numerical benchmarks (BBOB) and scalability to high dimensions (LSGO).
V-B1 Precision and Stability on Standard Benchmarks
Table I and Table II report the comparative results on BBOB-10D and BBOB-30D, respectively. Overall, L2E achieves consistently strong performance across a wide range of test instances:
  • Performance and Architectural Efficacy: L2E attains the lowest mean error on 31 out of 32 test instances. Specifically, L2E outperforms established learned baselines, such as GLHF and B2OPT, on ill-conditioned and multimodal functions (e.g., Rosenbrock_rot, Lunacek_bi_Ras). We attribute these gains to the stability-inducing unrolling design and the composite update structure, which promote controlled optimization dynamics in high-curvature regions. Furthermore, the Mamba-based variant (Ours) consistently improves upon the base model (Ours), particularly on non-separable landscapes (e.g., Ellipsoidal_hc). This finding suggests that the structured state-space operator provides an effective inductive bias for capturing population-wise variable couplings.
  • Trajectory and Stability Analysis: L2E demonstrates robust convergence behaviors, as visualized in Figure 4 and Figure 5. Specifically, the method exhibits faster early-stage progress with tighter confidence intervals. On unimodal functions (e.g., Bent Cigar), L2E quickly reaches a low-error regime, which indicates effective exploitation when the landscape is smooth. In contrast, on multimodal tasks (e.g., Gallagher 21Peaks), L2E maintains steady improvement without early stagnation. This sustained progress reflects the benefit of the Elite-Mask Fusion, which adaptively balances learned evolutionary proposals and proxy numerical guidance.
V-B2 Scalability to High-Dimensional Regimes
Results on the 1000D LSGO benchmark (Table III) further demonstrate the scalability of L2E. L2E ranks first across all six representative high-dimensional functions. Figure 6 (ECDF) and Fig. 7 (boxplots) illustrate the scalability of L2E in high-dimensional regimes. In contrast to existing learned optimizers, L2E exhibits consistent progress and reduced performance variance in the 1000D regime. This robustness stems from the structured operator design and the regularization strategies, i.e., bounded updates and soft-gated fusion. Consequently, these mechanisms promote controlled dynamics during optimization in high-dimensional spaces.
@@ V-B2 Scalability to High-Dimensional Regimes
Figure 5: Log-scaled convergence trajectories of our method and representative methods on the first 8 benchmark tasks from the BBOB-30D suite [20]. Complete results on all 16 functions are provided in the supplementary material (i.e., Fig S2).
@@
V-C Generalization to Real-World UAV Planning
We evaluate the robustness of L2E under distribution shifts using the UAV Path Planning benchmark. Although L2E is meta-trained exclusively on synthetic BBOB-30D functions, it is directly applied to 56 UAV planning instances (30D) without fine-tuning. As shown by the radar plots in Fig. 8, L2E achieves the best normalized performance across all seven terrain groups, with relative gains over the second-best baseline ranging from 3.4% to 14.2%. We attribute this empirical transferability to the stability-inducing unrolling formulation and the structured composite updates in L2E. Compared with policy-style learned optimizers that may become brittle when the constraint patterns and noise characteristics differ from training, L2E relies on structured population updates with stability-oriented regularization (e.g., bounded proposals and soft-gated numerical guidance). These design choices promote controlled dynamics under the strict evaluation budget, enabling more reliable progress on irregular UAV landscapes.
V-D Neuroevolution for Robot Control
@@
Finally, we evaluate the applicability of L2E to high-dimensional policy search in the Neuroevolution (NE) benchmark [22]. This setting introduces a distinct distribution shift from synthetic function optimization to control-oriented policy search. The benchmark includes six robotic tasks across two domains, ant and humanoid, where each domain comprises three complexity variants (ant-3/4/6 and humanoid-3/4/5). We compare L2E against GLEET [43], a specialized baseline that frames neuroevolution within an actor–critic reinforcement learning pipeline. This comparison tests whether a learned evolutionary optimizer can serve as a drop-in alternative to handcrafted population update heuristics in control scenarios.
Figure 9 reports normalized learning curves to enable a fair visual comparison across tasks with different reward scales. The normalization is defined as $y_{\text{plot}}=\frac{y-\text{mean}}{10^{\text{exp}}}.$ On representative mid-difficulty tasks (e.g., ant-4 and humanoid-4), L2E exhibits faster progress and improved final performance compared with the baseline. Moreover, the aggregated boxplots (Right) indicate that L2E achieves better terminal objective values with consistently reduced variance across all six tasks. These results suggest that the stability-inducing unrolling formulation and structured composite updates in L2E promote controlled policy search dynamics under noisy and high-variance feedback. Consequently, L2E improves robustness in neuroevolution-based control.
VI Ablation Analysis and Mechanistic Insights
To quantify the contribution of each component and motivate the design choices, we conduct systematic ablations on BBOB-10D. We focus on two questions: (1) how different parameterization strategies for the unrolled evolutionary operator affect optimization dynamics (shared vs. unshared weights across steps), and (2) how the core modules(i.e., ProxyGrad, SoftGate, and Mamba) influence performance and stability.
VI-A Shared vs. Unshared Unrolling Parameters
A pivotal design choice in neural unrolling determines whether to share the meta-parameters of the evolution operator $\mathcal{O}_{\mathrm{evo}}(\cdot;\bm{\omega})$ across unrolling steps $k=1,\dots,K$. This distinction establishes two primary configurations:
  • Shared ($\bm{\omega}_{\text{shared}}$): A single set of weights is utilized across all unrolled blocks, which enforces a time-invariant update rule.
  • Unshared ($\bm{\omega}^{k}$): Independent weights are assigned to distinct blocks, thereby enabling step-dependent behaviors that capture different phases of the optimization trajectory.